# Noncommutative Phantom BTZ Black Hole

### v1

@@ Abstract
This work explores the thermodynamic and geometric properties of phantom BTZ black holes within the framework of noncommutative spacetime, where noncommutative effects are incorporated via Lorentzian distributions for mass and charge. The resulting modifications in spacetime geometry introduce significant alterations to horizon structures and curvature singularities. A comprehensive and comparative thermodynamic analysis is conducted, examining the differences between phantom and ordinary matter cases. This includes an investigation of Hawking temperature, entropy, heat capacity, and stability criteria. Additionally, the black hole is analyzed as a thermodynamic heat engine, with its efficiency evaluated as a function of noncommutative parameters. Our findings highlight the profound impact of noncommutativity on the thermodynamic behavior and efficiency of phantom BTZ black holes, revealing new insights into the interplay between quantum spacetime effects and exotic field dynamics. The results indicate that noncommutative corrections not only modify the stability conditions of these black holes but also play a crucial role in governing phase transitions. Furthermore, we demonstrate that noncommutativity influences energy extraction processes, refining our understanding of black hole thermodynamics in lower-dimensional spacetimes and distinguishing the behavior of phantom and ordinary matter cases.
@@ 1 Introduction
The study of black holes has long captivated the scientific community, offering profound insights into the nature of spacetime, gravity, and quantum phenomena. These enigmatic objects, predicted by Einstein’s theory of general relativity [1], are not merely astrophysical entities but also thermodynamic systems, as first demonstrated by Bekenstein, Bardeen, Carter, and Hawking through their pioneering work on black hole entropy and radiation [2, 3, 4]. Their studies revealed that black holes obey laws analogous to the classical laws of thermodynamics, with entropy proportional to the event horizon area and temperature linked to Hawking radiation. This thermodynamic perspective has since provided profound insights into the quantum nature of spacetime, establishing black holes as pivotal in understanding gravity and quantum mechanics. Among the diverse black hole solutions, the Bañados-Teitelboim-Zanelli (BTZ) black hole in three-dimensional spacetime with a negative cosmological constant stands out as an elegant and analytically tractable model [5, 6, 7]. Despite its lower dimensionality, it retains essential features of higher-dimensional black holes, such as event horizons, Hawking radiation, and thermodynamic behavior. These attributes make the BTZ black hole a cornerstone for investigating the theoretical foundations of black hole thermodynamics and quantum gravity [8, 9, 10, 11, 12, 13, 14].
The pursuit of a unified theory of quantum gravity has led to profound modifications of classical spacetime concepts, particularly at the Planck scale where quantum and gravitational effects converge. A central feature of many quantum gravity frameworks, including string theory and loop quantum gravity, is the emergence of a minimal length scale, which imposes a fundamental limit on the resolution of spacetime. This idea is formalized in the Generalized Uncertainty Principle, which introduces quantum gravitational corrections to the standard uncertainty relations, predicting an inherent fuzziness in spacetime. Extensions such as the Extended Uncertainty Principle and the Extended Generalized Uncertainty Principle further incorporate large-scale curvature effects and unify short- and long-range corrections. While these principles have provided valuable insights into black hole physics, particularly for the BTZ black hole [15, 16, 17, 18, 19], noncommutative geometry offers a complementary perspective by encoding quantum gravitational effects directly into spacetime geometry [20]. Rooted in the idea that spacetime coordinates may no longer commute at microscopic scales, noncommutative geometry replaces point-like structures with smeared distributions, effectively regularizing singularities and introducing new corrections to classical solutions. This framework has found applications in string theory, quantum field theory, and cosmology, and has proven particularly insightful in black hole physics [21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. For BTZ black holes, noncommutative effects significantly modify their properties, such as altering the black hole horizon [31, 32, 33] and geodesic structure [34], affecting scattering phenomena [35], influencing thermodynamic behaviors [36, 37, 38, 39, 40], as well as introducing new effects like the modified Aharonov–Bohm effect [41]. The incorporation of smeared mass and charge distributions further underscores the critical role of noncommutative geometry in shaping the thermodynamic and dynamic properties of BTZ black holes [42].
The concept of black holes functioning as heat engines has garnered significant attention in recent years, offering a novel perspective on black hole thermodynamics [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. Johnson pioneered this idea by demonstrating the analogy between black holes and traditional heat engines, establishing a framework where thermodynamic cycles could be analyzed within the context of black hole physics [43, 44]. Subsequent studies extended this framework to various black hole solutions, such as charged black holes, rotating black holes, and those influenced by exotic fields, revealing intricate dependencies between thermodynamic variables and the efficiency of energy extraction [45, 51, 56]. In particular, the BTZ black hole has been investigated as a heat engine, providing insights into lower-dimensional thermodynamic cycles and the role of the cosmological constant as thermodynamic pressure [46]. The effects of dark energy [47], gravity’s rainbow [48], and massive gravity [50] have also been explored, demonstrating how these factors impact the efficiency and stability of black hole heat engines. Furthermore, novel configurations such as regular Bardeen black holes [57] and hairy black holes [58] have highlighted the diverse applications of this framework. Recent works continue to refine these analyses, including studies on modified Bardeen-AdS black holes [64], black holes in cavities [62], and accelerating black holes [55], showcasing the universality of black hole heat engines across different models and conditions.
On the other hand, the study of phantom fields has emerged as a significant area in black hole physics, providing deep insights into the interplay between exotic matter and spacetime geometry [69]. Phantom fields, characterized by violating the null energy condition, introduce negative energy densities and repulsive gravitational effects [70]. This behavior fundamentally alters the spacetime geometry, leading to scenarios such as horizon merging, shifts, or the formation of naked singularities. The Einstein-(anti)Maxwell-(anti)Dilaton theory provides a foundational framework for understanding these phenomena [71], wherein the coupling of scalar fields—either dilatonic or phantom—with Maxwell fields generates solutions exhibiting distinct thermodynamic and stability characteristics [72, 73]. Such fields have been extensively studied, with notable examples including the Gibbons-Rasheed solutions [74], their higher-dimensional generalizations, and related works exploring phantom contributions in various geometries. Despite concerns about potential quantum instabilities associated with phantom fields, subsequent analyses [75, 76] have suggested that such issues can be mitigated, enhancing their theoretical viability. Furthermore, cosmological evidence, such as the accelerated expansion of the universe driven by an exotic fluid with negative pressure that may possess a phantom nature [77, 78], highlights the importance of phantom fields [79]. These findings underscore their relevance in cosmological and black hole physics, inspiring methodologies for analyzing their thermodynamic properties and stability [80, 81, 82].
In the context of BTZ black holes, the inclusion of phantom energy introduces significant deviations from classical thermodynamic predictions [83]. Entropy no longer strictly adheres to the area law, while specific heat analysis reveals unconventional stability regimes. Furthermore, the interplay between phantom energy and phase transitions reshapes the thermodynamic landscape, with critical points governed by the unique properties of phantom fields. These modifications extend beyond classical physics, influencing quantum corrections such as Hawking radiation spectra and entanglement entropy [83]. Phantom BTZ black holes also provide a robust framework for exploring exotic states in dual field theories, highlighting their utility in understanding deviations from classical gravitational dynamics and enriching our understanding of quantum gravity in lower-dimensional spacetimes. Motivated by these advancements, this work investigates the thermodynamic and geometric properties of phantom BTZ black holes within the framework of noncommutative geometry. By introducing Lorentzian distributions for mass and charge, we derive a modified metric that encapsulates the effects of non-commutativity. This approach not only removes point-like singularities but also reveals new features in the spacetime structure, such as altered curvature profiles and horizon dynamics. Through a detailed thermodynamic analysis, we examine key quantities including the Hawking temperature, entropy, and heat capacity, elucidating the stability and phase transition behavior of these black holes. Furthermore, we analyze the black hole’s performance as a thermodynamic heat engine, highlighting the sensitivity of energy extraction efficiency to variations in noncommutative parameters. This study explores the intricate interplay between non-commutativity and phantom fields, aiming to shed light on the quantum aspects of gravitational systems.
This study is organized as follows: Section 2 delves into the geometry of the noncommutative phantom BTZ black hole. Section 3 provides a comprehensive thermodynamic analysis, including calculations of the Hawking temperature, entropy, and heat capacity, as well as an exploration of stability criteria. In Section 4, the black hole is treated as a heat engine, and its efficiency is analyzed, with particular emphasis on the role of noncommutative parameters in optimizing energy extraction. Finally, Section 5 summarizes our findings and discusses their potential implications for future research.
@@ 2 Phantom BTZ black holes in noncommutative space
$\Psi(r,\Theta)=c_{1}+\frac{\sqrt{\Theta}M}{\pi\sqrt{\Theta+r^{2}}}-\frac{\eta q^{2}}{8\pi^{2}}\left[\log\left(\sqrt{\Theta}+\sqrt{\Theta+r^{2}}\right)-\frac{1}{4}\log\left(\Theta+r^{2}\right)\right].$ (15)
Substituting Eq. (15) into Eq. (13), the metric function for the phantom BTZ black hole in noncommutative geometry is:
$f(r)=-M+\frac{\sqrt{\Theta}M}{\pi\sqrt{\Theta+r^{2}}}-\frac{\eta q^{2}}{8\pi^{2}}\left[\log\left(\sqrt{\Theta}+\sqrt{\Theta+r^{2}}\right)-\frac{1}{4}\log\left(\Theta+r^{2}\right)\right]-\Lambda r^{2}.$ (16)
@@
Figure 1: Variation of the metric function with respect to the event horizon radius $r_{H}$ for $\Lambda=-10^{-2}$, $M=0.1$ and $q=\frac{\pi}{2}$.
@@ 3 Thermodynamic Analysis
In this section, we investigate the thermodynamic properties of phantom BTZ black holes within the framework of noncommutative geometry. To achieve this, we express the black hole mass $M$ n terms of the event horizon radius $r_{H}$, the cosmological constant $\Lambda$, the charge $q$, and the noncommutative parameter $\Theta$, by imposing the condition $f\left(r_{H}\right)=0$. he mass is then given by the following expression:
@@
Figure 2: BTZ black hole mass $M$ as a function of the event horizon radius $r_{H}$ for $\eta=1$ and $q=\frac{\pi}{2}$. Panel (a) corresponds to $\Lambda=-10^{-2}$, and panel (b) corresponds to $\Lambda=-10^{-3}$.
@@
Figure 3: BTZ black hole mass $M$ as a function of the event horizon radius $r_{H}$ for $\eta=-1$ and $q=\frac{\pi}{2}$. Panel (a) corresponds to $\Lambda=-10^{-2}$, and panel (b) corresponds to $\Lambda=-10^{-3}$.
@@
By substituting the black hole mass (24) into the metric function (17) and evaluating Eq. (25), the Hawking temperature is calculated as:
@@
Figure 4: Hawking temperature $T$ as a function of the event horizon radius $r_{H}$ for $\eta=1$ and $q=\frac{\pi}{2}$. Panel (a) corresponds to $\Lambda=-10^{-2}$, and panel (b) corresponds to $\Lambda=-10^{-3}$.
@@
Figure 5: Hawking temperature $T$ as a function of the event horizon radius $r_{H}$ for $\eta=-1$ and $q=\frac{\pi}{2}$. Panel (a) corresponds to $\Lambda=-10^{-2}$, and panel (b) corresponds to $\Lambda=-10^{-3}$.
@@
$\mathcal{S}=\pi r_{H}+2\sqrt{\Theta}\log\Big{(}\frac{r_{H}}{r_{0}}\Big{)}-\frac{2\Theta}{\pi r_{H}},$ (28)
or equivalently,
$\mathcal{S}=S+2\sqrt{\Theta}\log\Big{(}\frac{S}{S_{0}}\Big{)}-\frac{2\Theta}{S}.$ (29)
where $S=\pi r_{H}$. This result demonstrates that the Bekenstein entropy is modified by the noncommutative parameter, while the entropy remains identical for both phantom and Maxwell black holes. Building on this, we turn our attention to another key thermodynamic quantity—the heat capacity—which provides critical insights into the stability and phase transitions of the black hole system. The heat capacity $C$ is defined as
$C=T\left(\frac{\partial\mathcal{S}}{\partial T}\right)_{q}.$ (30)
@@
Figure 6: Heat capacity $C$ as a function of the event horizon radius $r_{H}$ for $\eta=1$ and and $q=\frac{\pi}{2}$.
For the normal matter case, the heat capacity exhibits a divergence at a critical radius, marking the occurrence of a phase transition. For $r_{H}$ smaller than the critical radius, the heat capacity is negative, indicating thermodynamic instability. However, for horizon that are larger than the critical value, the heat capacity becomes positive, suggesting a stable thermodynamic regime. The magnitude of $C$ increases with $r_{H}$, reflecting the growing thermal capacity of the black hole as the event horizon expands. The effects of $\Lambda$ are apparent, with smaller values of $\Lambda$ as shown in panel (b), resulting in a gentler divergence and broader stability regimes compared to larger cosmological constant values, as given in panel (a).
@@
Figure 7: Heat capacity $C$ as a function of the event horizon radius $r_{H}$ for $\eta=-1$ and and $q=\frac{\pi}{2}$.
For the phantom matter case, the heat capacity displays more pronounced instabilities compared to the normal matter case. A divergence at the critical radius again signifies a phase transition. For small event horizons, the heat capacity is highly negative, reflecting enhanced thermodynamic instability induced by the phantom field. Beyond the critical radius, the heat capacity becomes positive, indicating the stabilization of the black hole for larger event horizons. The presence of phantom matter amplifies the divergence near the critical point, as evident in the steeper curves compared to the normal case. Additionally, as the cosmological constant decreases (panel (b)), the transition becomes more gradual, with a broader range of stability observed at larger horizons.
@@
Figure 8: Gibbs free energy $G$ as a function of the event horizon radius $r_{H}$ for $\eta=1$ and $q=\frac{\pi}{2}$.
@@
Figure 9: Gibbs free energy $G$ as a function of the event horizon radius $r_{H}$ for $\eta=-1$ and $q=\frac{3\pi}{2}$.
@@
$02\pi^{2}r_{c}^{2}-6\pi r_{c}\sqrt{\Theta}\left(2\pi-1+\log\Big{(}\frac{r_{c}}{r_{0}}\Big{)}\right)+\Theta\left(12\pi^{2}-30\pi+10-15\log\Big{(}\frac{r_{c}}{r_{0}}\Big{)}\right)=$ 0 . (44)
@@
0.3 0.390283 2.04167×$\times 10^{-2}$ 0.948005×$\times 10^{-2}$ 0.181 22
0.4 0.424061 1.96779×$\times 10^{-2}$ 0.869447×$\times 10^{-2}$ 0.187 37
0.5 0.452517 1.85648×$\times 10^{-2}$ 0.806325×$\times 10^{-2}$ 0.196 54
0.6 0.477323 1.73270×$\times 10^{-2}$ 0.754339×$\times 10^{-2}$ 0.207 80
0.7 0.499442 1.60607×$\times 10^{-2}$ 0.710545×$\times 10^{-2}$ 0.220 96
0.8 0.519487 1.48072×$\times 10^{-2}$ 0.672962×$\times 10^{-2}$ 0.236 10
0.9 0.537875 1.35850×$\times 10^{-2}$ 0.640218×$\times 10^{-2}$ 0.253 48
Table 1: Behavior of the critical parameters $r_{c}$, $T_{c}$, $p_{c}$ and $\frac{p_{c}r_{c}}{T_{c}}$ different values of the noncommutative parameter in the Maxwell case for and $q=\frac{\pi}{2}$
@@
0.3 3.31602 0.601952×$\times 10^{-2}$ 0.0730214×$\times 10^{-3}$ 0.0402 26
0.4 3.95094 0.507873×$\times 10^{-2}$ 0.0518237×$\times 10^{-3}$ 0.0403 16
0.5 4.52150 0.445444×$\times 10^{-2}$ 0.0397798×$\times 10^{-3}$ 0.0403 79
0.6 5.04534 0.400336×$\times 10^{-2}$ 0.0320773×$\times 10^{-3}$ 0.0404 26
0.7 5.53318 0.365876×$\times 10^{-2}$ 0.0267567×$\times 10^{-3}$ 0.0404 64
0.8 5.99211 0.338497×$\times 10^{-2}$ 0.0228761×$\times 10^{-3}$ 0.0404 96
0.9 6.42713 0.316096×$\times 10^{-2}$ 0.0199293×$\times 10^{-3}$ 0.0405 22
Table 2: Behavior of the critical parameters $r_{c}$, $T_{c}$, $p_{c}$ and $\frac{p_{c}r_{c}}{T_{c}}$ different values of the noncommutative parameter in the phantom case for and $q=\frac{\pi}{2}$.

### v2

# Thermodynamics and Heat Engine Behavior of Phantom BTZ Black Holes in Noncommutative Geometry

@@ Abstract
This study explores the thermodynamic and geometric properties of phantom BTZ black holes within a noncommutative spacetime framework, where noncommutativity is implemented through Lorentzian smearing of mass and charge distributions. The resulting metric exhibits significant modifications in curvature and horizon structure, particularly in the near-horizon regime. We perform a comparative thermodynamic analysis between phantom and Maxwell field cases, calculating quantities such as Hawking temperature, entropy, heat capacity, and Gibbs free energy. Our findings reveal that noncommutative corrections strongly affect phase transitions and stability conditions. Furthermore, we model the black hole as a heat engine and compute its efficiency, showing how noncommutative effects enhance or suppress energy extraction. This work underscores the interplay between spacetime fuzziness and exotic field dynamics in lower-dimensional gravity, offering new insights into quantum-modified black hole thermodynamics.
Keywords: Noncommutative Geometry, Phantom Fields, BTZ Black Holes, Black Hole Thermodynamics, Heat Engine Cycles,
PACS: 04.70.Dy, 11.10.Nx, 95.36.+x
@@ 1 Introduction
Black holes, once seen as purely theoretical solutions to Einstein’s field equations, are now recognized as critical arenas for probing the intersection of gravity, quantum mechanics, and thermodynamics. A pivotal shift occurred when Bekenstein and Hawking associated entropy with event horizon area and temperature with quantum radiation [1, 2, 3, 4], inaugurating black hole thermodynamics as a cornerstone of quantum gravity. In $(2+1)$ dimensions, the Bañados–Teitelboim–Zanelli (BTZ) black hole [5, 6, 7] provides an analytically tractable model that retains many essential features of higher-dimensional black holes, such as horizons, Hawking radiation, and thermodynamic behavior [8, 9, 10, 11, 12, 13, 14]. This has rendered the BTZ solution an effective laboratory for exploring lower-dimensional gravitational dynamics.
At Planckian scales, classical descriptions of spacetime cease to be adequate. Various quantum gravity frameworks—such as string theory and loop quantum gravity—propose a minimal length scale, leading to modifications of the Heisenberg uncertainty principle. Generalized, Extended, and Extended Generalized Uncertainty Principles (GUP, EUP, EGUP) incorporate these quantum corrections and have been fruitfully applied to BTZ thermodynamics [15, 16, 17, 18, 19]. An alternative route to incorporating short-distance effects lies in noncommutative (NC) geometry [20], wherein spacetime coordinates obey nontrivial commutation relations. This approach introduces a natural cutoff and regularizes classical singularities. In black hole physics, NC effects are typically modeled via smeared energy-momentum distributions, yielding modified metrics and altered thermodynamic profiles [21, 22, 23, 24, 25, 26, 27, 28]. In particular, NC corrections have been shown to deform the metric structure [29, 30, 31], alter geodesics [32], affect scattering amplitudes [33], and reshape entropy-area relations and phase transitions [34, 35, 36, 37, 38]. Additionally, they influence topological and quantum effects such as the Aharonov–Bohm phenomenon [39], gauge/gravity correspondence [40], and causal structure [41]. The Lorentzian-type smearing of matter and charge, in particular, yields improved regularity and refined thermodynamic behavior [43, 42]. These distributions are rooted in the notion of spatial fuzziness and local anisotropy—concepts that have long been central to models of self-gravitating systems [44].
Another compelling development in black hole physics involves phantom fields—exotic matter that violates the null energy condition and exhibits negative energy densities [46, 45]. Their presence can lead to nontrivial features such as horizon shifting, modified curvature profiles, and unconventional thermodynamic behavior [47, 48, 50, 49]. While concerns regarding quantum instability exist [51, 52], various studies have proposed stable, well-defined phantom configurations [55, 53, 54]. When coupled to BTZ spacetimes, phantom energy has been shown to influence entropy scaling, specific heat, and criticality [59, 56, 57, 58], broadening our understanding of black holes in lower dimensions.
Parallel to these developments, black holes have been reimagined as thermodynamic heat engines [60, 61, 62, 63, 65, 66, 67, 68, 70, 71, 72, 64, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 69]. Introduced by Johnson, this framework extends the standard thermodynamic analogy by treating the cosmological constant as a dynamical pressure and its conjugate as volume, enabling the construction of pressure–volume (PPVV) cycles analogous to those in classical heat engines. In this picture, black holes function as working substances that convert heat input into mechanical work, and their efficiency can be computed for specific thermodynamic trajectories (e.g., rectangular cycles in the PPVV plane). The BTZ black hole, due to its lower-dimensional simplicity and analytic tractability, has emerged as a compelling platform for studying such heat engines [63, 64]. Extensions have examined the influence of diverse physical phenomena on engine performance, including the effects of dark energy [65], gravity’s rainbow [66], massive gravity [68], and quantum corrections [69]. Modifications of the thermodynamic variables have been shown to affect the critical behavior, phase structure, and energy extraction efficiency. Moreover, nonstandard black hole configurations—such as Bardeen [75, 82], hairy [76], accelerating [73], and cavity-confined geometries [80]—have broadened the applicability of this paradigm. Recent works have also focused on benchmarking techniques [71], universal features in critical exponents [81], and thermodynamic geometry approaches to characterize engine performance and phase transitions [77, 79]. Even more exotic setups—including acoustic analogues [86], ModMax-AdS black holes [85], and magnetically charged configurations [84]—have confirmed the versatility of the heat engine formulation across a wide spectrum of gravitational systems. These efforts collectively demonstrate that the black hole heat engine is not merely a theoretical construct, but a powerful tool for probing the interplay between gravity, thermodynamics, and quantum corrections.
Motivated by these intersections, this study investigates BTZ black holes within noncommutative geometry in the presence of both phantom and Maxwell fields. Using Lorentzian-smeared energy-momentum tensors, we derive a modified metric and analyze its geometric and thermodynamic properties, including Hawking temperature, entropy, heat capacity, and Gibbs free energy. These quantities are examined to uncover phase transitions and assess the stability of the solutions. We further evaluate the black hole’s performance as a thermodynamic heat engine, quantifying the effects of noncommutativity and exotic matter content on energy conversion efficiency—an aspect not previously explored for these configurations. In fact, this work synthesizes and extends two prior directions in the literature: while Ref. [42] studied dynamic noncommutative BTZ black holes using both Gaussian and Lorentzian smearing functions—without coupling to exotic matter—and Ref. [59] analyzed BTZ black holes in the presence of both phantom and Maxwell fields—without invoking noncommutative geometry—our approach bridges these frameworks. By incorporating Lorentzian-type noncommutative corrections into BTZ black holes that include both phantom and Maxwell fields, and by conducting a comparative analysis of their heat engine behavior, we provide a unified platform for exploring the interplay between quantum spacetime structure and exotic matter in $(2+1)$-dimensional gravity.
This paper is structured as follows: Section 2 presents the construction of the noncommutative phantom BTZ metric. Section 3 details its thermodynamic properties and stability. Section 4 analyzes the black hole as a heat engine. Conclusions and outlook are provided in Section 5.
@@ 2 Phantom BTZ black holes in noncommutative space
$\Psi(r,\Theta)=c_{1}+\frac{\sqrt{\Theta}M}{\pi\sqrt{\Theta+r^{2}}}-\frac{\eta q^{2}}{8\pi^{2}}\left[\log\bigg(\frac{\sqrt{\Theta}+\sqrt{\Theta+r^{2}}}{r_{0}}\bigg)-\frac{1}{4}\log\bigg(\frac{\Theta+r^{2}}{r_{0}^{2}}\bigg)\right].$ (15)
The integration constant $c_{1}$ is determined by demanding that the metric reduces to the standard BTZ black hole form in the commutative limit $\Theta\to 0$, thereby identifying $c_{1}$ with the black hole mass parameter M-M. In addition, $r_{0}$ appears as an integration constant that serves as a reference length scale, ensuring that the arguments of the logarithmic terms remain dimensionless. Substituting Eq. (15) into Eq. (13), the metric function for the phantom BTZ black hole in noncommutative geometry is:
$f(r)=-M+\frac{\sqrt{\Theta}M}{\pi\sqrt{\Theta+r^{2}}}-\frac{\eta q^{2}}{8\pi^{2}}\left[\log\bigg(\frac{\sqrt{\Theta}+\sqrt{\Theta+r^{2}}}{r_{0}}\bigg)-\frac{1}{4}\log\bigg(\frac{\Theta+r^{2}}{r_{0}^{2}}\bigg)\right]-\Lambda r^{2}.$ (16)
@@
Figure 1: Variation of the metric function with respect to the event horizon radius $r_{H}$ for $\Lambda=-10^{-2}$, with $r_{0}=1$, $M=0.1$ and $q=\frac{\pi}{2}$.
@@ 3 Thermodynamic Analysis
In this section, we investigate the thermodynamic properties of both ordinary ($\eta=1$) and phantom ($\eta=-1$) BTZ black holes within the framework of noncommutative geometry. To achieve this, we express the black hole mass $M$ in terms of the event horizon radius $r_{H}$, the cosmological constant $\Lambda$, the charge $q$, and the noncommutative parameter $\Theta$, by imposing the condition $f(r_{H})=0$. The mass is then given by the following expression:
@@
Figure 2: BTZ black hole mass $M$ as a function of the event horizon radius $r_{H}$ for $\eta=1$, with $r_{0}=1$ and $q=\frac{\pi}{2}$. Panel (a) corresponds to $\Lambda=-10^{-2}$, and panel (b) corresponds to $\Lambda=-10^{-3}$.
@@
Figure 3: BTZ black hole mass $M$ as a function of the event horizon radius $r_{H}$ for $\eta=-1$, with $r_{0}=1$ and $q=\frac{\pi}{2}$. Panel (a) corresponds to $\Lambda=-10^{-2}$, and panel (b) corresponds to $\Lambda=-10^{-3}$.
@@
After evaluating the derivative at the horizon radius, the mass expression from Eq. (24) is substituted to obtain the Hawking temperature:
@@
Figure 4: Hawking temperature $T$ as a function of the event horizon radius $r_{H}$ for $\eta=1$, with $r_{0}=1$ and $q=\frac{\pi}{2}$. Panel (a) corresponds to $\Lambda=-10^{-2}$, and panel (b) corresponds to $\Lambda=-10^{-3}$.
@@
Figure 5: Hawking temperature $T$ as a function of the event horizon radius $r_{H}$ for $\eta=-1$, with $r_{0}=1$ and $q=\frac{\pi}{2}$. Panel (a) corresponds to $\Lambda=-10^{-2}$, and panel (b) corresponds to $\Lambda=-10^{-3}$.
@@
$S=\pi r_{H}+2\sqrt{\Theta}\log\Big(\frac{r_{H}}{r_{0}}\Big)-\frac{2\Theta}{\pi r_{H}}.$ (28)
This result demonstrates that the Bekenstein entropy is modified by the noncommutative parameter, while the entropy remains identical for both phantom and Maxwell black holes. Building on this, we turn our attention to another key thermodynamic quantity—the heat capacity—which provides critical insights into the stability and phase transitions of the black hole system. The heat capacity $C$ is defined as
$C=T\left(\frac{\partial S}{\partial T}\right)_{q}.$ (29)
@@
Figure 6: Heat capacity $C$ as a function of the event horizon radius $r_{H}$ for $\eta=1$, with $r_{0}=1$ and $q=\frac{\pi}{2}$.
For the normal matter case ($\eta=1$), the heat capacity is negative at small horizon radii, indicating local thermodynamic instability. As $r_{H}$ increases, $C$ crosses zero at a finite value of the horizon radius. This zero defines a critical horizon that separates the unstable branch ($C<0$) from the stable branch ($C>0$). Physically, at this point the system no longer exchanges heat with the surroundings, so the transition marks the onset of a thermodynamically stable regime. Beyond the critical horizon, $C$ remains positive and gradually tends toward zero as $r_{H}$ becomes large, reflecting the vanishing thermal response of very large black holes. The effect of the cosmological constant is illustrated by panels (a) and (b): larger $|\Lambda|$ displaces the critical horizon toward smaller values and broadens the stability region, while smaller $|\Lambda|$ leads to a larger critical horizon and a narrower stable domain.
@@
Figure 7: Heat capacity $C$ as a function of the event horizon radius $r_{H}$ for $\eta=-1$, with $r_{0}=1$ and $q=\frac{\pi}{2}$.
For the phantom matter case ($\eta=-1$), shown in Fig. 7, the heat capacity displays two distinct singularities. The first, located near $r_{H}\to 0$, is induced by the NC corrections and is common to both scenarios. The second singularity, however, appears only in the phantom case at a finite horizon radius and originates from the phantom contribution to the geometry. Between these two poles the heat capacity remains negative, signaling persistent instability. Only beyond the phantom–induced pole does $C$ become positive, marking the onset of a stable regime at sufficiently large horizons. As illustrated by panels (a) and (b), larger $|\Lambda|$ sharpen the singularities and broaden the stability domain, while smaller $|\Lambda|$ smooth them and reduce the extent of stability.
In summary, in both scenarios, the specific heat diverges very close to $r_{H}\to 0$ as a universal consequence of the noncommutative corrections. The subsequent behavior, however, differs markedly: in the normal case, stability is restored smoothly at the horizon where $C=0$, while in the phantom case, an additional singularity emerges at finite $r_{H}$, leading to a sharper and more abrupt instability. This contrast shows that phantom matter enriches the thermodynamic phase structure by introducing new singular features beyond the NC corrections.
@@
Figure 8: Gibbs free energy $G$ as a function of the event horizon radius $r_{H}$ for $\eta=1$, with $r_{0}=1$ and $q=\frac{\pi}{2}$.
@@
Figure 9: Gibbs free energy $G$ as a function of the event horizon radius $r_{H}$ for $\eta=-1$, with $r_{0}=1$ and $q=\frac{\pi}{2}$.
@@
$2\pi^{2}r_{c}^{2}-6\pi r_{c}\sqrt{\Theta}\left(2\pi-1+\log\Big(\frac{r_{c}}{r_{0}}\Big)\right)+\Theta\left(12\pi^{2}-30\pi+10-15\log\Big(\frac{r_{c}}{r_{0}}\Big)\right)=0.$ (43)
@@
[…]
Table 1: Behavior of the critical parameters $r_{c}$, $T_{c}$, $p_{c}$ and $\frac{p_{c}r_{c}}{T_{c}}$ for different values of the NC parameter in the Maxwell case with $r_{0}=1$ and $q=\frac{\pi}{2}$.
@@
[…]
Table 2: Behavior of the critical parameters $r_{c}$, $T_{c}$, $p_{c}$ and $\frac{p_{c}r_{c}}{T_{c}}$ for different values of the NC parameter in the phantom case with $r_{0}=1$ and $q=\frac{\pi}{2}$.